OpenStudy (anonymous):

Write y = x + 5 in standard form using integers.

jimthompson5910 (jim_thompson5910):

Standard form is Ax+By = C where A,B, C are integers Some books will make A > 0

OpenStudy (anonymous):

A. –x – 6y = 30 B. 6x – y = 30 C. –x + 6y = 30 D. –x + 6y = 5

jimthompson5910 (jim_thompson5910):

there has to be a typo in your original problem or one of the answer choices